OpenStudy (hba):

y varies directly as the cube of x. when x=0.5, y =2.5.if y=540 then x equals.

OpenStudy (hba):

@Praja_01

OpenStudy (hba):

@jim_thompson5910 Urgent help!

OpenStudy (hba):

@iNeedTheHelp

jimthompson5910 (jim_thompson5910):

y varies directly as the cube of x means y = kx^3

jimthompson5910 (jim_thompson5910):

when x=0.5, y =2.5 y = kx^3 2.5 = k*(0.5)^3 Solve for k

OpenStudy (hba):

Thanks a lot

jimthompson5910 (jim_thompson5910):

np

OpenStudy (hba):

@jim_thompson5910 k=2.5/0.025 < how to solve this without a calculator :D

OpenStudy (hba):

@jim_thompson5910 I got 100,what next ?

OpenStudy (hba):

@jim_thompson5910 ??

jimthompson5910 (jim_thompson5910):

you should have gotten k = 20

OpenStudy (hba):

I mean i got 20 what next

jimthompson5910 (jim_thompson5910):

y = kx^3 y = 20x^3 540 = 20x^3 Solve for x